OpenStudy (anonymous):

Determine whether the graphs of the pair of lines are parallel. x+8=y, y-x= -4

OpenStudy (anonymous):

Do you know about the slope of a line?

OpenStudy (anonymous):

No the next part is to find the slope

OpenStudy (anonymous):

okay, any linear equation in two variables represents a straight line on the cartesian plane.

OpenStudy (anonymous):

first step is to write the line's equation in the form of y = mx +C

OpenStudy (anonymous):

the first equation is already in this form. y = x + 8

OpenStudy (anonymous):

here m = 1 and C = 8

OpenStudy (anonymous):

any doubts?

OpenStudy (anonymous):

So for the second one is it y=x+4

OpenStudy (anonymous):

just move the x to the other side and it would become y = x - 4. the sign of four remains the same.

OpenStudy (anonymous):

here m = 1 and C = -4

OpenStudy (anonymous):

Is it clear?

OpenStudy (anonymous):

how do tell if they are parallel?

OpenStudy (anonymous):

the m represents the slope. Since both the lines have the same value of m (1 here ) they are parallel. This method is purely algebraic. you can also plot the graphs of the two given lines and check whether they are parallel.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

the slope of each line is 1. they are parallel.

OpenStudy (anonymous):

It was a pleasure